OpenStudy (anonymous):

Rhianna invests $1200 in stock and bonds. The stock pays 9% interest and the bonds pay 6% interest. If the total ANNUAL interest is $96, how much is invested in the stock alone?

OpenStudy (anonymous):

\[.09x + .06y = 9\] \[x + y = 1200\] \[x = -y + 1200\] \[.09(-y + 1200) + .06y = 96\] \[-.09y + 108 + .06y = 96\] \[-.03y = -12\] \[y = 400\] \[x + 400 = 1200\] \[x = 800\]